# A Generic Framework for Conformal Fairness

### v1

@@ 3.2 Conformal Fairness (CF) Theory
Filtering $\bm{\mathcal{D_{\mathrm{calib}}}}$: Group fairness metrics are evaluated on a subset of the population, defined by a condition on the data (i.e., membership in a group, true label value). For example, Demographic Parity is evaluated per group ($X\in g_{a}$ in definition), while Equal Opportunity is evaluated per group and true label ($Y=y,X\in g_{a}$ in definition). To formalize this notion, let $M$ denote a fairness metric (e.g. Equal Opportunity) and define $F_{M}:\mathcal{X}\times\mathcal{Y}\times\mathcal{G}\times\mathcal{Y}^{+}\to\{0,1\}$ be a filter function which maps a calibration point along with a group and positive label, $(\bm{x}_{i},y_{i},g,\tilde{y})$, to 00 or $1$ depending on whether the condition for the fairness metric, $M$, is satisfied. For Equal Opportunity, $F_{M}$ would instantiate to $F_{EO}(\bm{x}_{i},y_{i},g,\tilde{y}):=\bm{1}[\bm{x}_{i}\in g~{}\cap~{}y_{i}=\tilde{y}]$. We can filter $\mathcal{D_{\mathrm{calib}}}$ to be $\mathcal{D_{\mathrm{calib}}}_{(g,\tilde{y})}=\{(\bm{x}_{i},y_{i})\in\mathcal{D_{\mathrm{calib}}}~{}|~{}F_{M}(x_{i},y_{i},g,\tilde{y})=1\}$. By doing so, we provide guarantees regarding the conditional coverages as stated in Lemma 3.1
@@
Equation 1 holds if we replace $\{({\bm{x}}_{i},y_{i})\}$ with $\{({\bm{x}}_{i},\tilde{y})\}$ for a fixed $\tilde{y}\in{\mathcal{Y}}$.
@@ 3.3 Core Conformal Fairness (CF) Algorithm
2:     $\Lambda_{M}=\{\lambda\in\Lambda~{}|~{}\textsc{\text{Satisfy\_lambda}}(\mathcal{D_{\mathrm{calib}}},{\mathcal{Y}},{\mathcal{Y}}^{+},{\mathcal{G}},c,\lambda,F_{M})\}$triangleright Optimize for fairness
@@
7:procedure Satisfy_lambda($\mathcal{D_{\mathrm{calib}}}$, ${\mathcal{Y}}$, ${\mathcal{Y}}^{+}$,${\mathcal{G}}$, $c$,$\lambda$, $F_{M}$)
@@
[…]
@@
[…]
@@
17:         $\alpha_{\text{min}}=\text{min}(\text{label\_coverages}\text{[}(\cdot,\tilde{y})\text{]}-\text{interval\_widths}\text{[}(\cdot,\tilde{y})\text{]})$
@@
Using multiple $\bm{\lambda}$ thresholds: We also consider a classwise approach where we choose a $[\lambda^{0}_{opt},\ldots,\lambda^{k-1}_{opt}]=\bm{\lambda_{opt}}\in[0,1]^{K}$ for each of the $K$ classes. $\lambda^{i}_{opt}$ is only required to satisfy the closeness criterion for the $i^{\text{th}}$ class. One can achieve this by setting ${\mathcal{Y}}^{+}=\{\tilde{y}\}$ and repeating Lines 2 and 3 in Algorithm 1 for each $\tilde{y}\in{\mathcal{Y}}^{+}$. This allows for smaller $\lambda^{i}_{opt}$ to be chosen for most classes as they are no longer impacted by minority classes, which require a larger threshold to meet the closeness criterion.
@@ 3.4 Framework Extensibility
[…]
@@ 4.1 Setup
ACSIncome T $1,664,500$ ALL race$(9)$ $4$
ACSEducation T $1,664,500$ ALL race$(9)$ $6$
Name Type $(|{\mathcal{V}}|,|{\mathcal{E}}|)$ Labeled # Groups # Classes
Credit T/G $(30,000,~{}1,436,858)$ ALL age$(2)$ $4$
Pokec-n G $(66,569,~{}729,129)$ $8,797$ region$(2)$, gender$(2)$ $4$
Pokec-z G $(66,569,~{}729,129)$ $8,797$ region$(2)$, gender$(2)$ $4$
@@
Baseline: For each dataset and CP non-conformity score, we built a conformal predictor to achieve a coverage level of $0.91-\alpha=$ 0.9. Then, we assess fairness according to the specific fairness metric using the Satisfy_Lambda from Algorithm 1 for $\lambda=\hat{q}(\alpha)$.
@@ 4.2 Results
Figure 1: ACSIncome. The left two plots are efficiency results, while the right two are the fairness disparities for (a) APS, (b) RAPS, and (c) TPS. In all cases, our framework gives results at or better than the desired threshold and better than the baseline.
Figure 2: Credit. The top four plots are efficiency results, while the bottom four are the fairness disparities for (a) APS, (b) CFGNN, (c) DAPS, (d) RAPS, and (e) TPS. In all cases, our framework achieves the desired coverage gap better than the baseline, with a minor impact on efficiency.
@@
Figure 3: Pokec-n using both sensitive attributes. The top four plots are the efficiency results, while the bottom four are the fairness disparities for (a) APS, (b) CFGNN, (c) DAPS, (d) RAPS, and (e) TPS. CFGNN (b) and DAPS (c) achieve the desired fairness coverage thresholds better than standard CP methods.
@@
Table 4: ACSEducation. The worst-case fairness disparity, based on the Predictive Parity Proxy, with our method is below the desired $c$ threshold, while the average baseline disparity is much higher ($>0.30$) than all of the $c$ thresholds we consider.
@@ Appendix B Proofs
By computing $\hat{q}(\alpha)=\text{Quantile}(\frac{\lceil(n+1)(1-\alpha)\rceil}{n};\{s(\bm{x}_{i},\tilde{y})\}_{i=1}^{n})$, where $\tilde{y}$ is the fixed label. Using the assumption of exchangeability we have that $s(\bm{x}_{1},\tilde{y}),s(\bm{x}_{2},\tilde{y}),\ldots,s(\bm{x}_{n+1},\tilde{y})$ is an exchangeable sequence. Thus
@@ D.1 Datasets
In the fairness space, the American Community Services (ACS) datasets from the Folktables library are widely used (Ding et al., 2021). For ACSIncome, we used the standard ACSIncome dataset in Foltables however, we divided the targets into 4 classes by evenly dividing the income into 4 brackets. Race is the sensitive attribute and has 9 groups.
@@
Similar to ACSIncome, we used the ACS data and selected the Education Level as our target. We broke the education level into 6 groups {did not complete high school, has a high school diploma, has a GED, started an undergrad program, completed an undergrad program, and completed graduate or professional school}. ACSEducation also uses race as a sensitive attribute.
@@ D.2 Non-Conformity Scores
The most popular baseline when comparing CP method is APS (Romano et al., 2019). The scoring function works by sorting the softmax logits in descending order and accumulating the class probabilities until the correct class is included. For tighter prediction sets, randomization is introduced through a uniform random variable. Formally, if $\hat{\pi}({\bm{x}})_{(1)}\geq\hat{\pi}({\bm{x}})_{(2)}\geq\dots\geq\hat{\pi}({\bm{x}})_{(K-1)}$, $u\sim U(0,1)$, and $r_{y}$ is the rank of the correct label, then
@@ D.3 Hyperparameter Tuning
In the interest of reproducibility, the source code for the CF Framework is provided in the supplementary material.
@@
gamma uniform$(0,1)$
colsample_bytree uniform$(0.25,1.0)$
colsample_bylevel uniform$(0.25,1.0)$
colsample_bynode uniform$(0.25,1.0)$
subsample uniform$(0.5,1.0)$
@@ Appendix E Additional Results
Here we provide additional results and discourse for each dataset and experiment we discuss in the main paper.
@@ E.1 ACSEducation
Below we include results for the ACSEducation dataset, which is unique as it has a greater number of classes and is a custom dataset generated from the ACS datasets.
Figure E1: ACSEducation. Comparison of efficiencies when using the CF Framework without (top) and with (bottom) classwise lambdas. We observe that the efficiencies are better in the right plot. This is because $\forall_{i,\ldots,k}~{}\lambda_{\text{non-classwise}}\geq\lambda_{\text{classwise}}^{i}$ (k is the number of classes), which causes fewer labels to be included in the prediction set, thus improving efficiency with the classwise approach. For some experiments, the fairness disparity is 00 (e.g. APS and RAPS in the no-classwise setting), because the framework is producing the full prediction set–the trivial case–which means the coverage of $\tilde{y}\in\mathcal{Y}^{+}$ is $1.00$, thus causing the disparity to be 00.
Figure E2: ACSEducation. Comparison of fairness disparities when using the CF Framework without (top) and with (bottom) classwise lambdas. We observe that the fairness disparities are better in the left plot. This is because by using a single $\lambda$, only the hardest-to-satisfy label will be at or around the coverage gap, $c$, unlike classwise which ensures all labels will be at or around the coverage gap, $c$. Since fewer labels have coverages around the coverage gap, for non-classwise in (left) the likelihood of being above the threshold is limited - as opposed to the classwise approach (right).
@@ E.3.1 Pokec-n
Figure E6: Pokec-n. The efficiency (top) and fairness violation (bottom) plots when considering only gender as the sensitive attribute. Observe that the baseline disparity here is smaller than the baseline in intersectional fairness (see Figure 3). Thus, when controlling for $c=0.05$ coverage, there is a minimal change in efficiency across all the non-conformity scores.
Figure E7: Pokec-n. The efficiency (top) and fairness violation (bottom) plots when considering only region as the sensitive attribute. Compared to using gender in Figure E6, the prediction set sizes are larger for region when controlling for $c=0.05$, thus illustrating that exact performance will vary based on the sensitive attribute and group distribution.
@@ E.3.2 Pokec-z
Figure E8: Pokec-z. The efficiency (top) and fairness violation (bottom) plots when considering only gender as the sensitive attribute. Unlike Pokec-n - using gender (see Figure E6 - we find that the baseline is unfair for several values of c - exemplifying the auditing capabilities of the CF framework. This result also demonstrates how fairness can vary at different localities since Pokec-n and Pokec-z are disjoint subgraphs of the same graph.
Figure E9: Pokec-z. The efficiency (top) and fairness violation (bottom) plots when considering only region as the sensitive attribute. In these plots for $c\geq 0.1$, the baselines are fair for all score functions except for TPS. The baseline for TPS being ’unfair’ at $c=0.1$ suggests TPS sacrifices fairness for efficiency.
@@ E.6 Comparison with BatchGCP
To determine the value of $\lambda$, Jung et al. (2023) solve the convex optimization problem in 10, where $L_{q}$ is the pinball loss (Equation 11) - a function used to determine how close a threshold, $\hat{f}(x;\lambda)$, is to a specific quantile, $1-\alpha$, for a given score, $s$. Using the pinball loss, $\lambda^{*}$ the optimal $\lambda$ is computed and used to construct prediction sets.
[…]

### v2

@@ 3.2 Conformal Fairness (CF) Theory
[…]
@@
[…]
@@ 3.3 Core Conformal Fairness (CF) Algorithm
2:  $\Lambda_{M}=\{\lambda\in\Lambda~|~\textsc{\text{Satisfy\_lambda}}(\mathcal{D_{\mathrm{calib}}},{\mathcal{Y}},{\mathcal{Y}}^{+},{\mathcal{G}},c,\lambda,F_{M})\}$ $\triangleright$ Optimize for fairness
@@
8:  $\text{label\_coverages}=[0]_{({\mathcal{G}}_{i},y)\in{\mathcal{G}}\times{\mathcal{Y}}}$
9:  $\text{interval\_widths}=[0]_{({\mathcal{G}}_{i},y)\in{\mathcal{G}}\times{\mathcal{Y}}}$
@@
13:   $\text{interval\_widths}\text{[}(g,\tilde{y})\text{]}=\frac{1}{\lvert\mathcal{D_{\mathrm{calib}}}_{(g,\tilde{y})}\rvert+1}$ $\triangleright$ Uses Lemma 3.1
@@
17:   $)\alpha_{\text{min}}=\text{min}(\text{label\_coverages}\text{[}(\cdot,\tilde{y})\text{]}-\text{interval\_widths}\text{[}(\cdot,\tilde{y}$$\text{]})$
@@
Using multiple $\bm{\lambda}$ thresholds: We also consider a classwise approach where we choose a $\lambda^{0}_{opt},\ldots,\lambda^{k-1}_{opt}]=\bm{\lambda_{opt}}\in[0,1]^{K}$ for each of the $K$ classes. $\lambda^{i}_{opt}$ is only required to satisfy the closeness criterion for the $i^{\text{th}}$ class. One can achieve this by setting ${\mathcal{Y}}^{+}=\{\tilde{y}\}$ and repeating Lines 2 and 3 in Algorithm 1 for each $\tilde{y}\in{\mathcal{Y}}^{+}$. This allows for smaller $\lambda^{i}_{opt}$ to be chosen for most classes as they are no longer impacted by minority classes, which require a larger threshold to meet the closeness criterion.
@@ 3.4 Framework Extensibility
$\Pr[Y=\tilde{y}\mid\tilde{y}\in\mathcal{C}_{\lambda}(X),X\in g_{i}]=\underset{\text{Equal Opportunity over Demographic Parity}}{\underbrace{\frac{\Pr[\tilde{y}\in\mathcal{C}_{\lambda}(X)\mid Y=\tilde{y},X\in g_{i}]}{\Pr[\tilde{y}\in\mathcal{C}_{\lambda}(X)\mid X\in g_{i}]}}}~\cdot~\underset{\text{Conditional Label Probability}}{\underbrace{\Pr[Y=\tilde{y}~\mid~X\in g_{i}]}}$ (4)
@@ 4.1 Setup
ACSIncome T $1,664,500$ ALL race(9)(9) 44
ACSEducation T $1,664,500$ ALL race(9)(9) 66
Name Type $(|{\mathcal{V}}|,|{\mathcal{E}}|)$ # Labeled # Groups # Classes
Credit T/G (30,000,1,436,858)(30,000,~1,436,858) ALL age(2)(2) 44
Pokec-n G (66,569,729,129)(66,569,~729,129) $8,797$ region(2)(2), gender(2)(2) 44
Pokec-z G (66,569,729,129)(66,569,~729,129) $8,797$ region(2)(2), gender(2)(2) 44
@@
Baseline: For each dataset and CP non-conformity score, we built a conformal predictor to achieve a coverage level of $1-\alpha=0.9$. Then, we assess fairness according to the specific fairness metric using the Satisfy_Lambda from Algorithm 1 for $\lambda=\hat{q}(\alpha)$.
@@ 4.2 Results
Figure 1: ACSIncome. The top plots are efficiency results, while the bottom are the fairness disparities for (a) APS, (b) RAPS, and (c) TPS. In all cases, our framework gives results at or better than the desired threshold and better than the baseline.
Figure 2: Credit. The top plots are efficiency results, while the bottom are the fairness disparities for (a) APS, (b) CFGNN, (c) DAPS, (d) RAPS, and (e) TPS. In all cases, our framework achieves the desired coverage gap better than the baseline, with a minor impact on efficiency.
@@
Figure 3: Pokec-n using both sensitive attributes. The top plots are the efficiency results, while the bottom plots are the fairness disparities for (a) APS, (b) CFGNN, (c) DAPS, (d) RAPS, and (e) TPS. CFGNN (b) and DAPS (c) achieve the desired fairness coverage thresholds better than standard CP methods.
@@
Table 4: ACSEducation. The worst-case fairness disparity, based on the Predictive Parity Proxy, with our method is below the desired $c$ threshold, while the average baseline disparity is much higher (>0.30>0.30) than all of the $c$ thresholds we consider.
@@ Appendix B Proofs
[…]
@@ D.1 Datasets
In the fairness space, the American Community Services (ACS) datasets from the Folktables library are widely used (Ding et al., 2021). For ACSIncome, we used the standard ACSIncome dataset in Foltables; however, we divided the targets into 4 classes by evenly dividing the income into 4 brackets. Race is the sensitive attribute and has 9 groups.
@@
Similar to ACSIncome, we used the ACS data and selected the Education Level as our target. We broke the education level into 6 groups: {did not complete high school, has a high school diploma, has a GED, started an undergrad program, completed an undergrad program, and completed graduate or professional school}. ACSEducation also uses race as a sensitive attribute.
@@ D.2 Non-Conformity Scores
The most popular baseline when comparing CP method is APS (Romano et al., 2020). The scoring function works by sorting the softmax logits in descending order and accumulating the class probabilities until the correct class is included. For tighter prediction sets, randomization is introduced through a uniform random variable. Formally, if $\hat{\pi}({\bm{x}})_{(1)}\geq\hat{\pi}({\bm{x}})_{(2)}\geq\dots\geq\hat{\pi}({\bm{x}})_{(K-1)}$, $u\sim U(0,1)$, and $r_{y}$ is the rank of the correct label, then
@@ D.3 Hyperparameter Tuning
gamma uniform(0,1)(0,1)
colsample_bytree uniform(0.25,1.0)(0.25,1.0)
colsample_bylevel uniform(0.25,1.0)(0.25,1.0)
colsample_bynode uniform(0.25,1.0)(0.25,1.0)
subsample uniform(0.5,1.0)(0.5,1.0)
@@ Appendix E Additional Results
Here, we provide additional results and discourse for each dataset and experiment we discuss in the main paper.
@@ E.1 ACSEducation
Below, we include results for the ACSEducation dataset, which is unique as it has a greater number of classes and is a custom dataset generated from the ACS datasets.
Figure E1: ACSEducation. Comparison of efficiencies when using the CF Framework without (top) and with (bottom) classwise lambdas. We observe that the efficiencies are better in the right plot. This is because $\forall_{i,\ldots,k}~\lambda_{\text{non-classwise}}\geq\lambda_{\text{classwise}}^{i}$ (k is the number of classes), which causes fewer labels to be included in the prediction set, thus improving efficiency with the classwise approach. For some experiments, the fairness disparity is 0 (e.g., APS and RAPS in the no-classwise setting), because the framework is producing the full prediction set–the trivial case–which means the coverage of $\tilde{y}\in\mathcal{Y}^{+}$ is $1.00$, thus causing the disparity to be 0.
Figure E2: ACSEducation. Comparison of fairness disparities when using the CF Framework without (top) and with (bottom) classwise lambdas. We observe that the fairness disparities are better in the top plot. This is because by using a single $\lambda$, only the hardest-to-satisfy label will be at or around the coverage gap, $c$, unlike classwise, which ensures all labels will be at or around the coverage gap, $c$. Since fewer labels have coverages around the coverage gap, for non-classwise in (top), the likelihood of being above the threshold is limited - as opposed to the classwise approach (bottom).
@@ E.3.1 Pokec-n
Figure E6: Pokec-n. The efficiency (left) and fairness violation (right) plots when considering only gender as the sensitive attribute. Observe that the baseline disparity here is smaller than the baseline in intersectional fairness (see Figure 3). Thus, when controlling for $c=0.05$ coverage, there is a minimal change in efficiency across all the non-conformity scores.
Figure E7: Pokec-n. The efficiency (left) and fairness violation (right) plots when considering only region as the sensitive attribute. Compared to using gender in Figure E6, the prediction set sizes are larger for region when controlling for $c=0.05$, thus illustrating that exact performance will vary based on the sensitive attribute and group distribution.
@@ E.3.2 Pokec-z
Figure E8: Pokec-z. The efficiency (left) and fairness violation (right) plots when considering only gender as the sensitive attribute. Unlike Pokec-n - using gender (see Figure E6 - we find that the baseline is unfair for several values of c - exemplifying the auditing capabilities of the CF framework. This result also demonstrates how fairness can vary at different localities since Pokec-n and Pokec-z are disjoint subgraphs of the same graph.
Figure E9: Pokec-z. The efficiency (left) and fairness violation (right) plots when considering only region as the sensitive attribute. In these plots for $c\geq 0.1$, the baselines are fair for all score functions except for TPS. The baseline for TPS being ’unfair’ at $c=0.1$ suggests TPS sacrifices fairness for efficiency.
@@ E.6 Comparison with BatchGCP
To determine the value of $\lambda$, Jung et al. (2023) solve the convex optimization problem in 10, where $L_{q}$ is the pinball loss (Equation 11) - a function used to determine how close a threshold, $\hat{f}(x;\lambda)$, is to a specific quantile, $1-\alpha$, for a given score, $s$. Using the pinball loss, $\lambda^{*}$ - the optimal $\lambda$ - is computed and used to construct prediction sets.
$\lambda^{*}=\underset{\lambda}{\text{argmin}}~\mathbb{E}\big[L_{q}\big(\hat{f}(x;\lambda),s(x,y)\big)\big]$ (10)